LIU-PSB Meeting 194 (New RF bypasses update)
Participants: Simon Albright, Jean-Baptiste Bonnamy, Dario Carloni, Julie Coupard, Louis De Mallac, Joao Ferreira Bento, Alessandro Floriduz, Eve Fortescue-Beck, Georgi Minchev Georgiev, Greta Guidoboni, Klaus Hanke, Jan Hansen, Bettina Mikulec, Antony Newborough, Gabor Petrika, Serge Pittet, Tatiana Libera Rijoff, Andrea Santamaria Garcia, Jérôme Philippe Rodary.
Approval of Minutes
- The minutes of the LIU-PSB WG Meeting #193 were approved.
Communications
- There was a LIU Project meeting
on the 1st of June, minutes are not yet available. - A meeting
with SMB was held before this meeting:
- The alcove construction is ongoing and some work is foreseen for the Linac4 connection.
- SMB will communicate new WP holders.
- Report on the “Follow up on Bldg. 245” meeting held on the 31st of May 2017 (minutes):
- There is a surplus reserve of 942KCHF, which will be transferred to budget 67020. The spending profile will be adjusted accordingly.
- An updated schedule for building 245 is needed. The schedule should take into account all the modifications to be completed by November 17th so that other services can work with respect to that schedule.
Follow-up of Open Actions
A reminder needs to be sent for the following action:
- T. Dobers: "Report about the issues with the alignment of the PSB equipment and progress with the fabrication of a tilt measurement tool" (due on 2017-05-31).
A meeting concerning this topic is happening the 7th of June. The topic will be discussed again at the next meeting in two weeks.
- B. Holzer and J. Tan: "Provide specifications for the wire-scanners" (due on 2017-05-31).
It was agreed to start this action at the beginning of LS2:
- A. Newborough: "Write the ECR for the new dumping resistor installation" (due on 2017-06-30).
Bundle actions concerning the PSB-PS TL: advancing, no change in date.
- J-M. Lacroix: "Complete integration of PSB-PS transfer line.".
- B. Riffaud: "Provide proposal for the support of the equipment in the BTP line.".
- K. Hanke: "Organise discussion with ABT, RP, VSC and integration to clarify remaining open points for the BT/BTM/BTP transfer line integration (vacuum equipment, shielding, supports).".
- W. Bartmann: "Organise discussion with BI to finalize the beam instrumentation for the BT/BTM/BTP transfer lines for the post-LS2 situation.".
RF bypass measurements in PS with high-intensity beams by J. F. Bento
Tests were done in the PS on bypasses with 4 resistors, inserted in each corner of the clamp. The conclusion is that we can keep the present PSB bypasses during the post-LS2 era since the bypass could handle a peak current of 100 A during the test.
Comments:
- B. Mikulec asked how much time will it take to replace a clamp in case of failure.
- J. F. Bento answered that it can be done immediately. He remarked that the electric glue needed to connect the bypasses to the clamp needs to be applied at the moment of the installation anyways, to avoid losing its properties. Enough spares and material are available for a replacement in situ. He suggested talking to A. Blas for more details.
Tour de table
- S. Albright reported that a couple of MDs are scheduled this week to study phase jitter and emittance blow-up for LS2.
- G. Petrika mentioned an upcoming meeting on Thursday 8th of June to discuss the cooling needs of the upgraded PSB. K. Hanke reminded that the specifications of the cooling system have been documented and approved already. G. Petrika explained that the meeting will focus on optimizing the cooling system, not changing the specifications.
- J. Coupard reported on the ECR document concerning the SEM grid 4L1, which needs validation for the integration studies. She pointed out that the Linac4 operation and connection will be separated topics within the LIU. The budgets for the Linac4 connection will be discussed soon.
- J. P. Rodary reported on the safety in the CV local in building 245. He already provided the volume of water released in the CV local in case of leakage on the piping system. D. Carloni asked for an update of this number and will contact F. Boattini for that. He suggested exploring further the risk of leakage in other areas.
- D. Carloni reported on a meeting last week to discuss the access to the reference magnet room, located in the basement of building 245. He mentioned that the reference magnet is also located in the basement, close to the EN-CV equipment, which poses certain risks. He also mentioned that there are no oil retention tanks for the capacitors, which are in the next room, therefore the design of the draining system for B.245 shall be modified to avoid oil release towards the environment.
- G. M. Georgiev reported on the cabling. They are preparing the next YETS following the priorities for cabling.